OpenStudy (britanyashton):

When an object is not moving all the forces acting on it must sum to zero. The object is said to be in equilibrium. Two cables of different lengths hold a stoplight over an intersection. The force vectors being allied along the two cables are <20, 18> and <-20,12>. The magnitude of each vector is measured in pounds. A third force vector in this situation is the force due to gravity, and is straight downward. How much does the stoplight weigh?

OpenStudy (anonymous):

30, 18 + 12

OpenStudy (perl):

note that the drawing is not drawn to scale.

OpenStudy (perl):



OpenStudy (perl):

Because it is in equilibrum ΣF_x= 0 ΣF_y = 0 -20 + 20 + W_x = 0 W_x = 0 18 + 12 + W_y = 0 W_y = -30 so the weight is 30 units of force downward